OpenStudy (anonymous):

line that is perpendicular to -x+y=7 and passes through (-1,-1)

OpenStudy (anonymous):

First you have to find the slope of the line: -x + y = 7. The easiest way to do that is to change it to slope-intercept form. Can you do that?

OpenStudy (anonymous):

what is the eqution of the line

OpenStudy (anonymous):

do yo mean like y=mx+b

OpenStudy (anonymous):

I got y=-x-7

OpenStudy (anonymous):

Exactly, so what's the slope of THAT line?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Perfect. To be technical, let's say \[-\frac{ 1 }{ 1 }\] Now, do you know what you have to do to the slope of a line in order to get the slope of its perpendicular line? Two things...

OpenStudy (anonymous):

change to a positive and keep the y intercept?

OpenStudy (anonymous):

no right?

OpenStudy (anonymous):

Perpendicular lines have "opposite" and "reciprocal" slopes. So yes, it changes to a positive. And since it's 1/1, the reciprocal is still 1/1 (flipped). So the slope of the perpendicular line is positive 1. But to find the exact equation that passes through the point (-1, -1) with a slope of 1 (which we just determined), you have to use the point-slope formula. Do you know that? We can't just keep the old y-intercept.

OpenStudy (anonymous):

I don't remember point slope formula is it x2-x1 over y2-y1?

OpenStudy (anonymous):

No, that's almost the formula for slope (except the y's are on top). Point slope is this: \[y-y _{1}=m(x-x _{1})\] where \[(x _{1},y _{1})\] is a point on the line, and m is the slope. Can you plug in what you know? The regular x and y will stay x and y. Just replace the x1, y1, and m. Then we'll simplify.